# Continual Learning: Forget-free Winning Subnetworks for Video Representations

### v3

@@ 8 Conclusion
Inspired by Regularized Lottery Ticket Hypothesis (RLTH), which states that competitive subnetworks exist within a dense network in continual learning tasks, we introduce an interpretable continual learning approach referred to as Winning Subnetworks, WSN, which leverages re-used weights within dense networks in Task Incremental Learning (TIL) scenarios. We have also introduced variants of WSN, such as Soft-subnetwork (SoftNet), to address the overfitting in few-shot class incremental learning (FSCIL). Additionally, To overcome the limitation of WSN (sparse re-used weights) in video incremental learning (VIL), we have a newly proposed module as another variant of WSN that aims to find an adaptive and compact sub-module referred to as Fourier Subneural Operator (FSO) in Fourier space to encode videos in each video training session. The FSO finds reusable winning subnetworks in Fourier space, providing various bandwidths. We extend Fourier representations to various continual learning scenarios such as VIL, TIL, and FSCIL. Extensive experiments demonstrate the effectiveness of FSO in three continual learning scenarios. Overall, FSO’s representation markedly enhanced task performance at different convolutional representational levels, the higher layers for TIL and FSCIL and the lower layers for VIL.

### v4

@@ 9 Conclusion
Inspired by Regularized Lottery Ticket Hypothesis (RLTH), which states that competitive subnetworks exist within a dense network in continual learning tasks, we introduce an interpretable continual learning approach referred to as Winning Subnetworks, WSN, which leverages re-used weights within dense networks in Task Incremental Learning (TIL) and Task-agnostic Incremental Learning (TaIL) scenarios. We have also introduced variants of WSN, such as Soft-subnetwork (SoftNet), to address the overfitting in few-shot class incremental learning (FSCIL). Additionally, To overcome the limitation of WSN (sparse re-used weights) in video incremental learning (VIL), we have a newly proposed module as another variant of WSN that aims to find an adaptive and compact sub-module referred to as Fourier Subneural Operator (FSO) in Fourier space to encode videos in each video training session. The FSO finds reusable winning subnetworks in Fourier space, providing various bandwidths. We extend Fourier representations to various continual learning scenarios such as VIL, VaIL, TIL, and FSCIL. Extensive experiments demonstrate the effectiveness of FSO in three continual learning scenarios. Overall, FSO’s representation markedly enhanced task performance at different convolutional representational levels, the higher layers for TIL and FSCIL and the lower layers for VIL.